# Ahmed Hasanaath1, Aisha Alansari1, Ahmed Ashraf1, Chafik Salmane3, Hamzah Luqman1,2, Saad Ezzini1

### v1

@@ Abstract
Large language models (LLMs) have shown remarkable progress in reasoning abilities and general natural language processing (NLP) tasks, yet their performance on Arabic data, characterized by rich morphology, diverse dialects, and complex script, remains underexplored. This paper presents a comprehensive benchmarking study of multiple reasoning-focused LLMs, with a special emphasis on the newly introduced DeepSeek models, across a suite of fifteen Arabic NLP tasks. We experiment with various strategies, including zero-shot, few-shot, and fine-tuning. This allows us to systematically evaluate performance on datasets covering a range of applications to examine their capacity for linguistic reasoning under different levels of complexity. Our experiments reveal several key findings. First, carefully selecting just three in-context examples delivers an average uplift of over 13 F1 points on classification tasks—boosting sentiment analysis from 35.3 % to 87.5 % and paraphrase detection from 56.1 % to 87.0 %. Second, reasoning-focused DeepSeek architectures outperform a strong GPT o4-mini baseline by an average of 12 F1 points on complex inference tasks in the zero-shot setting. Third, LoRA-based fine-tuning yields up to an additional 8 points in F1 and BLEU compared to equivalent increases in model scale. The code is available at https://anonymous.4open.science/r/AraReasoner41299
@@ 1 Introduction
Arabic is spoken by more than 400 million people across 22 countries, making it one of the most widely used languages in the world. Despite its global significance, Arabic has traditionally been underrepresented in natural language processing (NLP) research [1]. This gap is partially due to the inherent complexity of the language, which features a rich morphological system, multiple dialects, and a non-Latin script. For instance, modern standard Arabic (MSA) differs considerably from various colloquial dialects in syntax and vocabulary, increasing the difficulty for models to generalize across tasks such as sentiment analysis, summarization, and translation. Moreover, diacritization and transliteration tasks pose unique challenges, as they require a nuanced understanding of phonological rules and script variations. These factors underscore the importance of dedicated research efforts to improve Arabic NLP capabilities.
@@
Large-scale language models (LLMs) have demonstrated outstanding performance across a wide range of NLP tasks [2]. Models from the GPT [3], Gemini [4], Deepseek [5], and LLaMA [6] families have shown exceptional capabilities in different NLP tasks. These models exhibit a strong understanding of natural language and express fluently in human language [7]. This performance can be attributed mainly to their model architecture and the pre-training on large-scale corpora. Recently, the newly introduced reasoning-based models, such as DeepSeek R1 and GPT-4o, have demonstrated impressive performance in several NLP tasks [8]. Although these models showed a good performance in English and other high-resource languages, their adaptability to Arabic remains an open question.
@@ 2 Related Work
Evaluating the performance of these LLMs requires robust, realistic, and task-specific benchmarks to ensure fair and meaningful comparisons. Due to the unique characteristics of the Arabic language, its complex morphology, rich diacritic system, and numerous dialects, Arabic-specific benchmarks have been developed to better capture these challenges. For instance, diacritization [9], the task of restoring diacritical marks to Arabic text, is a crucial task specific to Arabic. Transliteration [10] is another Arabic-specific task, often involving converting Arabizi (romanized Arabic script) into standard Arabic script. Additional tasks and benchmarks tailored to Arabic are discussed in more detail in Section 3.1. Prior studies have examined the performance of LLMs across multiple Arabic NLP tasks. For instance, [11] introduced recently ORCA, a challenging benchmark for Arabic language understanding, consisting of 60 different datasets across seven natural language understanding tasks. They used ORCA to offer a comprehensive comparison between 16 multilingual and Arabic pretrained language models. However, their evaluation is restricted to BERT-based models, focusing solely on understanding tasks and excluding generative capabilities and newer architectures such as decoder-only LLMs.
[12] assessed the performance of GPT-3.5 and GPT-4 across seven Arabic NLP tasks: sentiment analysis, translation, transliteration, paraphrasing, PoS tagging, summarization, and diacritization. Proving the outperformance of GPT-4 on six tasks. They also provide an evaluation pipeline that facilitates the evaluation of such models. However, their study is limited to just two closed-source models from the same family, leaving out a broader range of open-source and multilingual models. Abdelali et al. introduced LAraBench [13] for evaluating Arabic NLP and speech processing, comparing multiple models, including GPT-3.5, GPT-4, BLOOMZ, Jais-13b-chat, Whisper, and USM, using zero-shot and few-shot learning techniques to address 33 distinct tasks across 61 publicly available datasets. They found that state-of-the-art (SOTA) models generally outperformed the evaluated models in both zero-shot and few-shot settings, with a few exceptions in few-shot prompting.
Although prior studies such as ORCA, Taqyim, and LAraBench have significantly advanced the evaluation of Arabic NLP, there remains a notable gap in assessing the reasoning capabilities of more recent LLMs. Prior efforts have either focused on NLU using BERT-based models or a limited set of closed-source generative models. In this paper, we address this gap by introducing and evaluating a new set of SOTA models, DeepSeek-V3 [14], DeepSeek-R1 [14], GPT-o4, Qwen [15], QwQ [16] with a particular focus on their reasoning performance in Arabic NLP tasks.
@@ 3 Benchmark
Figure 2 shows the pipeline of AraReasoner. Fifteen Arabic NLP tasks were selected to evaluate the reasoning capabilities of different reasoning-based LLMs on these tasks. A prompt instruction selection process was followed for each task to identify the most effective prompt (see Appendix C Prompt Selection). Subsequently, extensive evaluation experiments were conducted under various prompting strategies, including zero-shot and in-context learning. In addition, we compared the reasoning-based models not only against each other but also against non-reasoning LLMs to highlight the impact of reasoning abilities on task performance.
@@ 3.1 Tasks and Datasets
Sentence classification. This category encompasses a diverse range of sentence-level classification NLP tasks. Eight tasks have been involved in this work: sentiment analysis (SA), dialect detection (DD), sarcasm detection (SarD), hate speech detection (HSD), offensive speech detection (OSD), claim detection (CD), Stance detection (StD), and Paraphrase detection (PD). The SA task involves determining a sentence’s emotional tone or opinion, usually as positive, negative, or neutral. We use the AJGT Corpus, which comprises tweets in the Jordanian dialect [17], and we used the split proposed by [18]. The DD task distinguishes MSA from dialectal Arabic, whereas the SarD task identifies whether a given tweet conveys sarcasm. For DD and SarD tasks, we utilize the ArSarcasm corpus [19]. The HSD is the task of identifying language that expresses hatred, discrimination, or hostility toward individuals or groups based on traits like race, gender, religion, or ethnicity. The OSD task captures rude or disrespectful language that may insult but is not necessarily individual or group-targeted. For HSD and OSD tasks, we employ the dataset proposed by [20]. The CD task involves checking the factuality of a claim. The StD task detects the writer’s stance towards a particular target or topic. The goal is to detect whether the writer agrees, disagrees, discusses, or has no opinion on a given topic. For CD and StD tasks, we used the dataset presented in [21] as it provides labels for both tasks. PD checks whether a pair of sentences conveys the same meaning, and we use the dataset introduced by [22] for this task, which consists of Arabic sentence pairs annotated for semantic equivalence.
@@
Natural language generation. This category includes five principal tasks in NLG. It involves summarization (SUM), paraphrasing (PAR), transliteration (TRL), machine translation (MT), and generative question answering (GQA) tasks. The SUM task involves generating a concise and coherent version of a longer text while preserving its key information. For this task, we use the AraSum [23] as a training set and EASC as a testing set [24]. This approach was chosen because AraSum provides clear and concise summaries, which enhance fine-tuning quality. The PAR task focuses on rephrasing a given text while maintaining its original meaning. For this task, we utilize the Arabic Paraphrased Parallel Synthetic dataset (APPSD) dataset [25], excluding Arabic Paraphrasing Benchmark (APB) [26], which we use for testing. The TRL task involves converting text from one script to another while preserving its pronunciation. It enables non-Arabic speakers to approximate the pronunciation of Arabic words by representing them using Latin letters. We use the BOLT dataset [27] for this task. We also evaluate the reasoning-based models on the MT task using the Arabic-English split of the United Nations Parallel Corpus v1.0 (UNv1) [28]. Moreover, we evaluate these models on the generative QA task that involves generating natural language answers to questions based on a given context. For this task, we utilize the XTREME benchmark [29].
@@
Linguistic analysis. This category encompasses POS tagging and (WSD) tasks. The PoS tagging task (PoS) assigns a grammatical category (e.g., noun, verb, adjective) to each word in a sentence. We use the full Universal Dependencies v2.3 corpus [30] for PoS tagging evaluation experiments. The WSD task identifies the correct meaning of a word with multiple possible senses based on its context. We utilize the Arabic WSD benchmark [31] for this task.
@@ 3.2.2 DeepSeek
We evaluated the reasoning-based models of the DeepSeek [32], particularly DeepSeek-R1 and its distilled Qwen-based variants. These models aim to advance reasoning in language models through reinforcement learning (RL) and effective distillation.
DeepSeek-R1-Zero is trained purely via large-scale RL without supervised fine-tuning. This model demonstrated emergent reasoning behaviors such as self-verification and reflection [33]. However, it struggles with repetition, readability, and language mixing.
DeepSeek-R1 addresses the limitations of DeepSeek-R1-Zero by introducing cold-start supervised fine-tuning before RL. This resulted in improving stability and performance across math, code, and reasoning tasks [33].
DeepSeek-V3 DeepSeek-V3 is a 671B-parameter mixture-of-experts (MoE) model with multi-head latent attention and DeepSeekMoE for efficient training. It uses an auxiliary-loss-free multi-token prediction and load balancing for better performance. Trained on 14.8 trillion tokens with supervised fine-tuning and reinforcement learning, it matches top models with stable training [34].
Qwen Based Distilled DeepSeek Models We focus on the distilled variants of DeepSeek-R1 that are based on the Qwen2.5 series, including 1.5B, 7B, 14B, and 32B models. These models transfer reasoning capabilities from DeepSeek-R1 through a distillation process that preserves core performance characteristics while significantly reducing model size. Among them, DeepSeek-R1-Q32B has demonstrated competitive performance, reportedly outperforming OpenAI’s o1-mini model on several standard benchmarks [33]. However, it is important to note that their effectiveness in Arabic NLP tasks remains unexplored.
@@ 3.2.3 Qwen
QwQ-32B [16] is a 32B parameter reasoning model from the Qwen series, developed to tackle complex tasks via advanced RL techniques. This model showed strong reasoning capabilities, achieving competitive results on math, coding, and general benchmarks [16]. Furthermore, it performed on par with much larger models such as DeepSeek-R1 (671B, 37B active) and OpenAI-o1-mini.
@@ 4.2 Experimental Setup
In our experiments, we accessed the LLMs via APIs and conducted all fine-tuning locally. For fine-tuning, we focused on select models from the DeepSeek family. Appendix B Prompt Strategies lists the prompt strategies followed in this work, and Appendix D Implementation Details gives more information on inference settings, fine-tuning configurations, and hardware specifications.
@@ 4.2.1 Data Split
The datasets were divided into Train, Dev, and Test splits. The Dev set was used for prompt selection. Train was used for fine-tuning, while Test served for final evaluation. See Appendix A Data statistics for more dataset details.
@@ 4.2.3 Prompts Selection
We started the evaluation process by selecting the appropriate prompt for each task using the DeepSeek R1-Q14B model. We designed six prompt variants for each task by crossing three styles with two languages (English and Arabic) as shown in Appendix C Prompt Selection. The three prompt styles can be classified as Instructive, Interrogative, and Role-playing. Instructive prompts use an imperative tone, Interrogative prompts take the form of a question, and Role-playing prompts present the instruction as if the model is performing a task in a specific role.
Table 1 presents the performance of the DeepSeek R1-Q14B model in a zero-shot learning setup, using three prompt types written in both English and Arabic. The results highlight the importance of tailoring prompt styles to each task. As shown in the table, the best performance for most tasks, particularly in SC and LA tasks, was achieved using Arabic instructive prompts. For NLG and some LA tasks, English prompts, especially instructive ones, outperformed the others. Additionally, the lowest performance across nearly all tasks, except for the TRL and PoS tagging tasks, was observed when using role-playing prompts, regardless of language. Based on these findings, all subsequent experiments were conducted using the best-performing prompt for each task.
@@ 4.3 Results
This section outlines the performance evaluations of the selected LLMs across a variety of Arabic NLP tasks. The findings are detailed in several tables. Firstly, Table 2 presents the performance data for 15 tasks, comparing leading reasoning models such as R1-671B, QWQ-32B, and GPT 04-mini, under various prompting strategies. Following this, Table 3 illustrates the performance variations across different model sizes within the DeepSeek R1 family for the chosen tasks, also employing diverse prompting strategies. Furthermore, Table 4 offers a comparative analysis of the fine-tuning performance of different R1 model sizes on the selected tasks. Finally, Table 5 contrasts the performance of non-reasoning models, specifically Deepseek V3-685B and GPT-4o, against the reasoning R1-671B model in a zero-shot setting. These tables collectively provide a detailed view of how model choice, size, and prompting techniques influence outcomes on the benchmarked Arabic NLP tasks. Visual representations of these results can be found in Appendix E Evaluation Results.
Table 2: Reasoning models capabilities across multiple tasks. DS: Deepseek; ZS: Zero-shot; 3S: 3 shots; 5S: 5 shots.
@@
Table 5: Reasoning vs Non-reasoning models in zero-shot mode.

### v4

# Hamzah Luqman1,2, Saad Ezzini1 1

@@ ABSTRACT
Large language models (LLMs) have shown remarkable progress in reasoning abilities and general natural language processing (NLP) tasks, yet their performance on Arabic data, characterized by rich morphology, diverse dialects, and complex script, remains underexplored. This paper presents a comprehensive benchmarking study of multiple reasoning-focused LLMs, with a special emphasis on the newly introduced DeepSeek models, across a suite of fifteen Arabic NLP tasks. We experiment with various strategies, including zero-shot, few-shot, and fine-tuning. This allows us to systematically evaluate performance on datasets covering a range of applications to examine their capacity for linguistic reasoning under different levels of complexity. Our experiments reveal several key findings. First, carefully selecting just three in-context examples delivers an average uplift of over 13 F1 points on classification tasks—boosting sentiment analysis from 35.3 % to 87.5 % and paraphrase detection from 56.1 % to 87.0 %. Second, reasoning-focused DeepSeek architectures outperform a strong GPT o4-mini baseline by an average of 12 F1 points on complex inference tasks in the zero-shot setting. Third, LoRA-based fine-tuning yields up to an additional 8 points in F1 and BLEU compared to equivalent increases in model scale. The code is available at Project Repository
@@ 1 Introduction
Arabic is spoken by more than 400 million people across 22 countries, making it one of the most widely used languages in the world. Despite its global significance, Arabic has traditionally been underrepresented in natural language processing (NLP) research alturayeif2022mawqif. This gap is partially due to the inherent complexity of the language, which features a rich morphological system, multiple dialects, and a non-Latin script. For instance, modern standard Arabic (MSA) differs considerably from various colloquial dialects in syntax and vocabulary, increasing the difficulty for models to generalize across tasks such as sentiment analysis, summarization, and translation. Moreover, diacritization and transliteration tasks pose unique challenges, as they require a nuanced understanding of phonological rules and script variations. These factors underscore the importance of dedicated research efforts to improve Arabic NLP capabilities.
@@
Large-scale language models (LLMs) have demonstrated outstanding performance across a wide range of NLP tasks chang2024survey. Models from the GPT radford2018gpt, Gemini team2023gemini, Deepseek bi2024deepseek, and LLaMA touvron2023llama families have shown exceptional capabilities in different NLP tasks. These models exhibit a strong understanding of natural language and express fluently in human language alyafeai2306taqyim. This performance can be attributed mainly to their model architecture and the pre-training on large-scale corpora. Recently, the newly introduced reasoning-based models, such as DeepSeek R1 and GPT-4o, have demonstrated impressive performance in several NLP tasks cheng2025empowering. Although these models showed a good performance in English and other high-resource languages, their adaptability to Arabic remains an open question as arabic remains a morphologically rich yet low-resource language antoun2020arabert.
@@ 2 Related Work
Evaluating the performance of these LLMs requires robust, realistic, and task-specific benchmarks to ensure fair and meaningful comparisons. Due to the unique characteristics of the Arabic language, its complex morphology, rich diacritic system, and numerous dialects, Arabic-specific benchmarks have been developed to better capture these challenges. For instance, diacritization alkhamissi2020diacretization, the task of restoring diacritical marks to Arabic text, is a crucial task specific to Arabic. Transliteration shazal2020transliteration is another Arabic-specific task, often involving converting Arabizi (romanized Arabic script) into standard Arabic script. Additional tasks and benchmarks tailored to Arabic are discussed in more detail in Section 3.1. Prior studies have examined the performance of LLMs across multiple Arabic NLP tasks. For instance, elmadany2022orca introduced recently ORCA, a challenging benchmark for Arabic language understanding, consisting of 60 different datasets across seven natural language understanding tasks. They used ORCA to offer a comprehensive comparison between 16 multilingual and Arabic pretrained language models. However, their evaluation is restricted to BERT-based models, focusing solely on understanding tasks and excluding generative capabilities and newer architectures such as decoder-only LLMs.
Taqyimitw alyafeai2023taqyim assessed the performance of GPT-3.5 and GPT-4 across seven Arabic NLP tasks: sentiment analysis, translation, transliteration, paraphrasing, PoS tagging, summarization, and diacritization, proving the outperformance of GPT-4 on six tasks. They also provide an evaluation pipeline that facilitates the evaluation of such models. However, their study is limited to just two closed-source models from the same family, leaving out a broader range of open-source and multilingual models. Abdelali et al. introduced LAraBench abdelali2023larabench for evaluating Arabic NLP and speech processing, comparing multiple models, including GPT-3.5, GPT-4, BLOOMZ, Jais-13b-chat, Whisper, and USM, using zero-shot and few-shot learning techniques to address 33 distinct tasks across 61 publicly available datasets. They found that state-of-the-art (SOTA) models generally outperformed the evaluated models in both zero-shot and few-shot settings, with a few exceptions in few-shot prompting.
Although prior studies such as ORCA, Taqyim, and LAraBench have significantly advanced the evaluation of Arabic NLP, there remains a notable gap in assessing the reasoning capabilities of more recent LLMs. Prior efforts have either focused on NLU using BERT-based models or a limited set of closed-source generative models. In this paper, we address this gap by evaluating a new set of SOTA reasoning-oriented models including DeepSeek-V3 liu2024deepseek_v3, DeepSeek-R1 liu2024deepseek_v3, GPT-4o, Qwen bai2023qwen, and QwQ qwq32b on a diverse suite of Arabic NLP tasks.
@@ 3 Benchmark
Figure 2 shows the pipeline of AraReasoner. We selected fifteen Arabic NLP tasks to evaluate how well reasoning-based LLMs perform across a diverse range of language understanding challenges. A prompt instruction selection process was followed for each task to identify the most effective prompt (see Appendix C). Subsequently, extensive evaluation experiments were conducted under various prompting strategies, including zero-shot and in-context learning. In addition, we compared the reasoning-based models not only against each other but also against non-reasoning LLMs to highlight the impact of reasoning abilities on task performance.
@@ 3.1 Tasks and Datasets
Sentence classification. This category encompasses a diverse range of sentence-level classification NLP tasks. Eight tasks have been involved in this work: sentiment analysis (SA), dialect detection (DD), sarcasm detection (SarD), hate speech detection (HSD), offensive speech detection (OSD), claim detection (CD), stance detection (StD), and paraphrase detection (PD). The SA task involves determining a sentence’s emotional tone or opinion, usually as positive, negative, or neutral. We use the AJGT Corpus, which comprises tweets in the Jordanian dialect alomari2017arabic, and we used the split proposed by nagoudi-etal-2022-arat5. The DD task distinguishes MSA from dialectal Arabic, whereas the SarD task identifies whether a given tweet conveys sarcasm. For both DD and SarD tasks, we utilize the ArSarcasm corpus farha2020arabic, which includes annotated labels for both sarcasm and dialect. The HSD is the task of identifying language that expresses hatred, discrimination, or hostility toward individuals or groups based on traits like race, gender, religion, or ethnicity. The OSD task captures rude or disrespectful language that may insult but is not necessarily individual or group-targeted. For HSD and OSD tasks, we employ the dataset proposed by mubarak2020overview. The CD task involves checking the factuality of a claim. The StD task detects the writer’s stance towards a particular target or topic. The goal is to detect whether the writer agrees, disagrees, discusses, or has no opinion on a given topic. For CD and StD tasks, we used the dataset presented in khouja2020stance as it provides labels for both tasks. PD checks whether a pair of sentences conveys the same meaning, and we use the dataset introduced by seelawi2019nsurl for this task, which consists of Arabic sentence pairs annotated for semantic equivalence.
@@
Natural language generation. This category includes five principal tasks in NLG. It involves summarization (SUM), paraphrasing (PAR), transliteration (TRL), machine translation (MT), and generative question answering (GQA) tasks. The SUM task involves generating a concise and coherent version of a longer text while preserving its key information. For this task, we use the AraSum kahla-etal-2021-cross as a training set and EASC as a testing set el2010using. This approach was chosen because AraSum provides clear and concise summaries, which enhance fine-tuning quality. The PAR task focuses on rephrasing a given text while maintaining its original meaning. For this task, we utilize the Arabic Paraphrased Parallel Synthetic dataset (APPSD) dataset al2024arabic, excluding Arabic Paraphrasing Benchmark (APB) alian2019towards, which we use for testing. The TRL task involves converting text from one script to another while preserving its pronunciation. It enables non-Arabic speakers to approximate the pronunciation of Arabic words by representing them using Latin letters. We use the BOLT dataset shazal2020unified for this task. We also evaluate the reasoning-based models on the MT task using the Arabic-English split of the United Nations Parallel Corpus v1.0 (UNv1) ziemski2016united. Moreover, we evaluate these models on the generative QA task that involves generating natural language answers to questions based on a given context. For this task, we utilize the XTREME benchmark siddhant2020xtreme.
@@
Linguistic analysis. This category encompasses POS tagging and (WSD) tasks. The PoS tagging task (PoS) assigns a grammatical category (e.g., noun, verb, adjective) to each word in a sentence. We use the full Universal Dependencies v2.3 corpus nivre2018universal for PoS tagging evaluation experiments. The WSD task identifies the correct meaning of a word with multiple possible senses based on its context. We utilize the Arabic WSD benchmark el2021arabic for this task.
@@ 3.2.2 DeepSeek
We evaluated the reasoning-based models of the DeepSeek guo2025deepseek, particularly DeepSeek-R1 and its distilled Qwen-based variants. These models aim to advance reasoning in language models through reinforcement learning (RL) and effective distillation.
DeepSeek-R1-Zero is trained purely via large-scale RL without supervised fine-tuning. This model demonstrated emergent reasoning behaviors such as self-verification and reflection guo2025deepseek_r1. However, it struggles with repetition, readability, and language mixing.
DeepSeek-R1 addresses the limitations of DeepSeek-R1-Zero by introducing cold-start supervised fine-tuning before RL. This resulted in improving stability and performance across math, code, and reasoning tasks guo2025deepseek_r1.
DeepSeek-V3 DeepSeek-V3 is a 671B-parameter mixture-of-experts (MoE) model with multi-head latent attention and DeepSeekMoE for efficient training. It uses an auxiliary-loss-free multi-token prediction and load balancing for better performance. Trained on 14.8 trillion tokens with supervised fine-tuning and reinforcement learning, it matches top models with stable training liu2024deepseek.
Qwen Based Distilled DeepSeek Models We focus on the distilled variants of DeepSeek-R1 that are based on the Qwen2.5 series, including 1.5B, 7B, 14B, and 32B models. These models transfer reasoning capabilities from DeepSeek-R1 through a distillation process that preserves core performance characteristics while significantly reducing model size. Among them, DeepSeek-R1-Q32B has demonstrated competitive performance, reportedly outperforming OpenAI’s o1-mini model on several standard benchmarks guo2025deepseek_r1. However, it is important to note that their effectiveness in Arabic NLP tasks remains unexplored.
@@ 3.2.3 Qwen
QwQ-32B qwq32b is a 32B parameter reasoning model from the Qwen series, developed to tackle complex tasks via advanced RL techniques. This model showed strong reasoning capabilities, achieving competitive results on math, coding, and general benchmarks qwq32b. Furthermore, it performed on par with much larger models such as DeepSeek-R1 (671B, 37B active) and OpenAI-o1-mini.
@@ 4.2 Experimental Setup
In our experiments, we accessed the LLMs via APIs and conducted all fine-tuning locally. For fine-tuning, we focused on select models from the DeepSeek family. Appendix B lists the prompt strategies followed in this work, and Appendix D gives more information on inference settings, fine-tuning configurations, and hardware specifications.
@@ 4.2.1 Data Split
The datasets were divided into Train, Dev, and Test splits. The Dev set was used for prompt selection. Train was used for fine-tuning, while Test served for final evaluation. See Appendix A for more dataset details.
@@ 4.2.3 Prompts Selection
We started the evaluation process by selecting the appropriate prompt for each task using the DeepSeek R1-Q14B model. We designed six prompt variants for each task by crossing three styles with two languages (English and Arabic) as shown in Appendix C. The three prompt styles can be classified as Instructive, Interrogative, and Role-playing. Instructive prompts use an imperative tone, Interrogative prompts take the form of a question, and Role-playing prompts present the instruction as if the model is performing a task in a specific role.
Table 1 presents the performance of the DeepSeek R1-Q14B model in a zero-shot learning setup, using three prompt types written in both English and Arabic. The results highlight the importance of tailoring prompt styles to each task. As shown in the table, the best performance for most tasks, particularly in SC and LA tasks, was achieved using Arabic instructive prompts. For NLG and some LA tasks, English prompts, especially instructive ones, outperformed the others. Additionally, the lowest performance across nearly all tasks, except for the TRL and PoS tagging tasks, was observed when using role-playing prompts, regardless of language Based on these findings, all subsequent experiments were conducted using the best-performing prompt types(INS, INT, RP) for each task, combined with different few-shot configurations (zero-shot, 3-shot, and 5-shot) to assess the impact of few-shot prompting on task performance.
@@ 4.3 Results
This section outlines the performance evaluations of the selected LLMs across a variety of Arabic NLP tasks. The findings are detailed in several tables. Firstly, Table 2 presents the performance data for 15 tasks, comparing leading reasoning models such as R1-671B, QWQ-32B, and GPT 04-mini, under various prompting strategies. Following this, Table 3 illustrates the performance variations across different model sizes within the DeepSeek R1 family for the chosen tasks, also employing diverse prompting strategies. Furthermore, Table 4 offers a comparative analysis of the fine-tuning performance of different R1 model sizes on the selected tasks. Finally, Table 5 contrasts the performance of non-reasoning models, specifically Deepseek V3-685B and GPT-4o, against the reasoning R1-671B model in a zero-shot setting. These tables collectively provide a detailed view of how model choice, size, and prompting techniques influence outcomes on the benchmarked Arabic NLP tasks. Visual representations of these results can be found in Appendix E.
Table 2: Reasoning models across tasks. DS: Deepseek; ZS: zero-shot; 3S: 3 shots; 5S: 5 shots.
@@
AVG 73.31 82.94 91.13
Table 5: Reasoning vs. non-reasoning models in zero-shot mode.